### v1

# GRPose: Learning Graph Relations for Human Image Generation with Pose Priors

Abstract
Recent methods using diffusion models have made significant progress in human image generation with various additional controls such as pose priors. However, existing approaches still struggle to generate high-quality images with consistent pose alignment, resulting in unsatisfactory outputs. In this paper, we propose a framework delving into the graph relations of pose priors to provide control information for human image generation. The main idea is to establish a graph topological structure between the pose priors and latent representation of diffusion models to capture the intrinsic associations between different pose parts. A Progressive Graph Integrator (PGI) is designed to learn the spatial relationships of the pose priors with the graph structure, adopting a hierarchical strategy within an Adapter to gradually propagate information across different pose parts. A pose perception loss is further introduced based on a pretrained pose estimation network to minimize the pose differences. Extensive qualitative and quantitative experiments conducted on the Human-Art and LAION-Human datasets demonstrate that our model achieves superior performance, with a 9.98% increase in pose average precision compared to the latest benchmark model. The code is released on *******.
Figure 1: GRPose generates results by better aligning with pose prior and scaling outputs to $512\times 512$ pixels. The first row compares it with ControlNet while the second row visualizes the pose alignment across different base models. ControlNet exhibits pose alignment in certain body parts such as arms and incorrect actions like handshakes.
**footnotetext: Xun Yang is the corresponding author.
Introduction
Human image generation aims to synthesize high-quality images under specific conditions based on a series of prompts, such as canny edge, pose and depth (Yang et al. 2023). Its diverse applications range from animation (Corona et al. 2024) and game production (Pan et al. 2024) to other fields. Early methods (Men et al. 2020; Ma et al. 2017) primarily adopted variational autoencoders (VAEs) (Kingma and Welling 2013) or Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), leveraging a source image to synthesize target images with specific human attributes. While these methods achieve control through the reference appearance, the synthesis process is unstable and the training heavily depends on the distribution of the source images. Recently, Stable Diffusion (SD) (Rombach et al. 2022) and its variants (Podell et al. 2023) have been developed to address these limitations, in which high-fidelity human images are synthesized with the help of a prompt. Considering the data availability and computation costs, controllable diffusion models (Zhang, Rao, and Agrawala 2023; Mou et al. 2024; Zhao et al. 2024) further introduce a learnable control branch into the frozen SD model, enabling spatial control of the generative results by providing conditions such as depth maps and segmentation masks.
ControlNet (Zhang, Rao, and Agrawala 2023) has emerged as a new benchmark by fine-tuning the frozen SD model with trainable copy parameters, enabling spatial control through conditional inputs. HumanSD (Ju et al. 2023b) adopts a heat-guided loss to achieve pose control. While these methods synthesize images with satisfactory semantics and style, they struggle to produce high-quality outputs based on given pose priors, often resulting in unrealistic body alignment (as shown in Figure 1). A key challenge in existing approaches is the poor alignment with the given pose priors. We observed that these methods incorporate human pose information into frozen generative models based on Euclidean space, which inadequately represents the nonlinear and higher order relationships between different parts of pose priors, particularly in terms of joint connections and overall coordination.
In this study, we propose to capture the topological relationships among different pose parts by leveraging a graph structure to extract intrinsic associations. We propose a framework named Graph Relation Pose (GRPose) to guide Stable Diffusion. We follow the ControlNet utilizing both a text prompt and a pose prior to generate human images, and primarily fine-tune the Adapter which is a set of trainable copy parameters. We aim to establish a graph topological structure on pose priors and latent representations, adopting Graph Convolutional Networks (GCNs) (Kipf and Welling 2016) to effectively capture the higher order relationships between different pose parts. We introduce a graph construction mechanism called the Progressive Graph Integrator (PGI). PGI treats each spatial part of pose priors and latent representations as graph nodes, mapping the spatial relationships of the pose into a graph structure. This mechanism distills pose information into each layer within the Adapter through a hierarchical structure, allowing gradual refinement of the pose while preserving the authenticity of the appearance. Additionally, we propose a novel pose perception loss, which adopts a pre-trained pose estimation network to minimize the pose differences between outputs and original images, further encouraging alignment of the synthesized output with pose priors.
Our contributions are summarized as follows:
  • We propose a framework called Graph Relation Pose (GRPose), leveraging a graph structure to more comprehensively capture the topological and spatial information of pose priors for human image generation, offering a new perspective in the field.
  • We design a Progressive Graph Integrator (PGI) to explicitly capture the intrinsic associations between different pose parts, distilling pose information into each layer within the Adapter.
  • We introduce a novel pose perception loss that utilizes a pre-trained pose estimation network to minimize pose differences.
  • We conduct extensive qualitative and quantitative experiments on the Human-Art (Ju et al. 2023a) and LAION-Human datasets (Ju et al. 2023c). Our GRPose achieves superior performance compared to advanced methods across multiple evaluation metrics, particularly in terms of pose guidance alignment.
Figure 2: Overview of Graph Relation Pose (GRPose). It consists of Diffusion Pipeline, Graph Pose Adapter and Pose Perception Loss. The Pose Encoder uses a coupling structure in alongside the Progressive Graph Integrator (PGI) to capture graph relationships between different pose parts. The Pose Perception Loss adopts a pre-trained pose estimation network to regularize the pose alignment in the feature space.
Related Work
Pose-Guided Human Image Generation
Traditional pose-guided human generation methods take a specific source image and pose condition as input to generate images that retain the appearance of the source image while presenting a specified pose. Some studies (Lv et al. 2021; Ma et al. 2021; Yang et al. 2021) are based on GANs or VAEs to convert the task as a conditional generation. Bhunia et al.  (Bhunia et al. 2023) leveraged a Person Image Diffusion Model (PIDM) to synthesize images by learning a noise distribution and proposed a cross-attention-based Texture Diffusion Module (TDB) to align the relationship between appearance and pose information. Zhang et al.  (Zhang et al. 2022) proposed a dual-task pose transformer and introduced an auxiliary task, connecting different branches to obtain correlation between features by building attitude conversion modules. Shen et al.  (Shen et al. 2023) introduced a Progressive Conditional Diffusion Model (PCDM), which incrementally bridges the gap between character images under target poses and source poses through a three-stage process. Despite the significant progress of existing methods, challenges remain in handling complex poses. Our study seeks to achieve more realistic human images by introducing graph associations and a pose perception loss.
Controllable T2I Diffusion Model
Large Diffusion models (Rombach et al. 2022; Ramesh et al. 2022) can generate high-quality images under a set of prompts. Ho et al.  (Ho and Salimans 2022) proposed classifier-free guidance that combines conditional prediction with unconditional prediction. ControlNet (Zhang, Rao, and Agrawala 2023) introduces a trainable control branch by incorporating a weight copy of the diffusion model, allowing spatial control of generative images using depth maps, segmentation masks and more. HumanSD (Ju et al. 2023b) finetunes the diffusion model with a pose-guided heatmap loss. T2I-Adapter (Mou et al. 2024) trains a lightweight adapter to guide the frozen SD model. UniControlNet (Zhao et al. 2024) utilizes two additional control modules to achieve control over multiple conditions. Although these methods demonstrate strong control capabilities, they struggle with handling complex pose relationships. To address this, we introduce graph structures to capture high-level associations in the pose-guided generation process.
Method
Overview of GRPose
Our aim is to generate high-quality human images conditioned on pose priors. Our proposed GRPose consists of three main components: Diffusion Pipeline, Graph Pose Adapter and Pose Perception Loss, as shown in Figure 2. Given a pose condition $c_{p}\in\mathbb{R}^{H\times W\times C}$ and a text prompt as inputs, the CLIP text encoder (Radford et al. 2021) converts the text into its embedding $c_{t}$. In the Pose-Guided Diffusion Pipeline, we use Stable Diffusion (Rombach et al. 2022) as the base diffusion model, which includes a VAE for image encoding and decoding, and a U-Net for noise estimation. Specifically, an image is fed into the encoder of the VAE to obtain the latent representation $z_{0}$. The optimization objective is to maximize the logarithmic likelihood of the data distribution, as defined:
$\mathcal{L}_{d}=E_{\mathbf{z}_{t},t,\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon_{t}-\epsilon_{\theta}\left(\mathbf{z}_{t},\mathbf{c}_{t},\mathbf{c}_{p},t\right)\right\|^{2}\right],$ (1)
where $z_{t}$ represents the latent representation at time step $t$, $\mathcal{N}$ represents the standard normal distribution, $\epsilon_{\theta}$ represents the noise prediction network and $\epsilon_{t}$ denotes real noise.
Within the entire framework, the Graph Pose Adapter is a trainable component that encodes the pose condition $c_{p}$ into a graph structure and integrates it into the Adapter through a hierarchical structure. At the beginning of each encoder layer in the Adapter, the encoded pose and the current latent representation are fed into the Progressive Graph Integrator (PGI) to capture the topological relationships between different pose parts through graph learning. This process fine-tunes the Adapter to convey control signals to the SD model, producing the synthesized image $\hat{x}$. Additionally, to further encourage alignment of the synthesized output with pose priors, the pose perception loss is formulated using a pre-trained pose estimation network to quantify the pose differences between outputs and the original images.
Progressive Graph Integrator
We adopt a pose encoder with $L$ layers ($L$=4, similar to the Adpater) to capture multi-level scales of pose priors within a hierarchical structure, where a PGI is incorporated into each encoding layer and its corresponding layer in the Adapter. In the PGI, each spatial point in the encoded pose prior $x_{p}$ and the latent representation $x_{l}\in\mathbb{R}^{H\times W\times C}$ is treated as a node for constructing respective graph structures, as shown in Figure 3. By doing this, both local features of spatial relationships in the pose prior and latent representation are captured, where each point in $x_{p}$ or $x_{l}$ represents a field-of-view of $(2^{L-1}\times 2^{L-1})$ in the original images. Then, we proceed to construct the feature graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the node set is defined as $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ with $N=H\times W$ representing the number of nodes. The graph edges between nodes $v_{i}$ and $v_{j}$ are established by applying a K-nearest neighbors ($K$NN) search algorithm:
$N_{K}(v_{i})=\{v_{j}\in V\mid d(v_{i},v_{j})\leq K\},$ (2)
where $d$ represents the $L_{2}$ distance between two node vectors. Notably, we incorporate positional encoding $p_{i}$ into each node ( $v_{i}\leftarrow v_{i}+p_{i}$), where $p_{i}$ denotes the spatial positions of the corresponding patch. A set of edge $\mathcal{E}$ can be obtained as:
=$\mathcal{E}=\{(v_{i},v_{j})|v_{j}\in N_{K}(v_{i}),i\neq j\},$ (3)
Then, two separate graph structures, $\mathcal{G}_{p}$ and $\mathcal{G}_{l}$, are formed to represent the pose prior $x_{p}$ and latent representastion $x_{l}$ respectively.
With these two graphs, we employ graph convolution (GC) layers to facilitate the association of features between nodes within each graph. The GC layer is divided into two steps: aggregation and update. In the aggregation step, features of neighboring nodes are gathered as described:
$F_{agg}=\sum_{j\in\mathcal{K}(i)}\left(\mathcal{A}_{ij}\cdot X_{j}\right),$ (4)
where $\mathcal{K}(i)$ is the neighboring node $i$ determined by the $K$NN algorithm, $\mathcal{A}\in\mathbb{R}^{N\times N}$ is an adjacency matrix, and $X$ denotes the features of nodes. In the update step, each node integrates the aggregation features to update its current representation. This process is expressed as:
$X^{\prime}_{i}=\boldsymbol{\phi}\left(\boldsymbol{\Theta}\cdot X_{i}\oplus(1-\boldsymbol{\Theta})\cdot F_{agg}\right),$ (5)
where $\boldsymbol{\phi}$ denotes the activation function, $\boldsymbol{\Theta}$ is a learnable parameter. To balance the degrees of different nodes, the adjacency matrix $\mathcal{A}$ is normalized as:
$\hat{\mathcal{A}}=D^{-\frac{1}{2}}(\mathcal{A}+I)D^{-\frac{1}{2}},$ (6)
where $D$ denotes the diagonal degree of $\mathcal{A}$, $I$ denotes the identity matrix. The GC process can be described as:
$\hat{X}=\boldsymbol{\phi}\left(\hat{\mathcal{A}}XW\right)+X,$ (7)
where $W$ denotes the weight matrix. The GC layer distills information across different parts for both pose prior and latent representation. Take the pose prior as the example, this process is summarized as:
$\hat{X_{p}}=\mathrm{GC}(\mathcal{G}_{p},X_{p}),$ (8)
We further fuse the graph features of pose and latent through a fuse layer, which captures the cross-modal interactions between them. The fuse layer consists of three convolution blocks. The fused features are then used to construct a graph $\mathcal{G}_{fuse}$, which is fed into an additional GC layer to capture the complex associations to refine the feature representation.
Figure 3: Details of Progressive Graph Integrator (PGI). The pose prior $x_{p}$ and latent representation $x_{l}$ are gridded to construct graphs $\mathcal{G}_{p}$ and $\mathcal{G}_{l}$ respectively, where GCNs are employed to fuse and update the information.
Pose Perception Loss
We introduce a novel Pose Perception Loss to enhance the pose alignment during the image synthesis process under the guidance of a pose perception network. Our approach employs a pre-trained pose estimation network, specifically designed to accurately identify and comprehend the poses of individuals within images. By encoding both the generated image $\hat{x}$ and the original image $x$, the discrepancies between them can be captured effectively, the it can be defined as:
$\mathcal{L}_{p}=\frac{1}{hw}\sum_{i=0}^{h}\sum_{j=0}^{w}\left\|\varphi_{p}\left(\hat{x}_{ij}\right)-\varphi_{p}\left(x_{ij}\right)\right\|_{2}^{2},$ (9)
where $\varphi_{p}$ denotes the encoder of the pose estimation network, which extract the features from the pose backbone, $h$ and $w$ denotes the width and height of both $x$ and $\hat{x}$ respectively and $\left\|\cdot\right\|_{2}$ signifies the Euclidean distance. By directly minimizing the pose differences in the feature space, our method is capable of generating images that are not only visually more realistic but also consistent with pose priors. This loss function can be integrated as a plug-and-play component into existing image generation frameworks, providing additional pose guidance to these models.
The optimization target of the training process is:
$\mathcal{L}=\mathcal{L}_{d}+\alpha*\mathcal{L}_{p}.$ (10)
where $\mathcal{L}_{d}$ represents the diffusion loss, $\alpha$ respresents the weight of the pose perception loss $\mathcal{L}_{p}$. The parameters of the Adapter are updated through this loss function.
Table 1: Results on Human-Art and LAION-Human datasets. The best results and the second best results are marked in green and blue respectively. Results marked with asterisk (*) are evaluated on the released models.
Dataset Methods AP (%) $\uparrow$ SAP (%) $\uparrow$ PCE $\downarrow$ FID $\downarrow$ KID $\downarrow$ CLIP-Score $\uparrow$
Human-Art SD* (Rombach et al. 2022) 0.24 55.71 2.30 11.53 3.36 33.33
T2I-Adapter (Mou et al. 2024) 27.22 65.65 1.75 11.92 2.73 33.27
ControlNet (Zhang, Rao, and Agrawala 2023) 39.52 69.19 1.54 11.01 2.23 32.65
Uni-ControlNet (Zhao et al. 2024) 41.94 69.32 1.48 14.63 2.30 32.51
HumanSD (Ju et al. 2023b) 44.57 69.68 1.37 10.03 2.70 32.24
GRPose (Ours) 49.50 70.84 1.43 13.76 2.53 32.31
LAION-Human SD* (Rombach et al. 2022) 0.73 44.47 2.45 4.53 4.80 32.32
T2I-Adapter* (Mou et al. 2024) 36.65 63.64 1.62 6.77 5.44 32.30
ControlNet* (Zhang, Rao, and Agrawala 2023) 44.90 66.74 1.55 7.53 6.53 32.31
Uni-ControlNet (Zhao et al. 2024) 50.83 66.16 1.41 6.82 4.52 32.39
HumanSD (Ju et al. 2023b) 50.95 65.84 1.25 5.62 7.48 30.85
GRPose (Ours) 57.01 67.20 1.29 6.52 4.65 32.12
Experiments
Experimental Settings
Datasets. We evaluated our model on the Human-Art (Ju et al. 2023a) and LAION-Human (Ju et al. 2023c) datasets. The Human-Art dataset comprises 50,000 high-quality images from 5 real-world and 15 virtual scenarios, featuring human bounding boxes, key points and textual descriptions. The LAION-Human dataset consists of approximately 1 million image-caption pairs, filtered by high image quality and human estimation confidence scores, using a diverse range of human activities and more realistic images. On LAION-Human dataset we randomly selected 200,000 samples for training and 20,000 samples for testing.
Implementation Details. For a fair comparison, our diffusion pipeline adopts the same Stable Diffusion 1.5111https://huggingface.co/runwayml/stable-diffusion-v1-5 as previous methods. We used the Adam optimizer (Kingma and Ba 2014) with an initial learning rate of $1\times 10^{-5}$. We trained models on 8×8$\times$8 × NVIDIA L40S-48GB, using the Pytorch and PytorchLightning [^2] https://lightning.ai/docs/pytorch/stable/. The batch size was set as 6, and the number of epochs was 20. The weight of pose loss $\alpha$ was set as 0.01 and we adopted a gradual constraint strategy, adding pose perception loss in the last 5 epochs. We replaced the text prompts with an empty string with a probability of 0.5. For the PGI, we set the parameter $K$ of the KNN algorithm to 9. During inference, the number of steps $T$ for the DDIM sampler was set to 50. We employed mmpose [^3] https://github.com/open-mmlab/mmpose framework as the pretrained pose estimation networks. It is worth noting that, in the evaluation phase we used a different pre-trained pose estimation network during inference process to assess the accuracy of the pose in the generated images. This ensures the independence and reliability of the evaluation.
Metrics. We adopted Average Precision (AP), Similarity Average Precision (SAP) and Person Count Error (PCE) (Cheong, Mustafa, and Gilbert 2022). Higher scores in AP and SAP indicate better pose alignment. Additionally, we employed the Fréchet Inception Distance (FID) (Heusel et al. 2017) and Kernel Inception Distance (KID) (Bińkowski et al. 2018) to assess the quality of the generated images, and used the CLIP-Score (Radford et al. 2021) to evaluate the alignment between the image and text.
Figure 4: Results of Multi-Pose Generation. Our model outperforms ControlNet in generating multiple poses.
Comparison with SOTA Methods
Quantitative results comparing our method with other state-of-the-art (SOTA) approaches are shown in Table 1. The experimental results indicate that our GRPose achieved the highest AP and SAP. On the Human-Art dataset, our model attained AP and SAP of 49.50% and 70.84% respectively, bringing an improvement of 9.98% in AP compared to ControlNet. Similarly, on the LAION-Human dataset, our model showed a consistent improvement of 6.06% in AP compared to HumanSD (Ju et al. 2023b). Notably, our method significantly improved pose alignment accuracy, as evidenced by the AP, SAP, and PCE metrics. We observed that SD1.5 using only a text prompt without pose conditions lacked the ability for pose alignment. Although controllable diffusion models such as ControlNet and T2I-Adapter used pose conditions for spatial control, their capabilities of pose alignment were limited due to incorporating it without considering the intrinsic high-order associations between different pose parts. While retaining pose alignment, our model also maintained good performance in image quality and text alignment with minor decline in scores.
Figure 5: Comparison of different base diffusion models for our GRPose. We compared SD1.5, Anime Art and Realistic models of different styles. GRPose effectively enhances each model with superior pose alignment.
We further compared the visualizations with other methods (Zhang, Rao, and Agrawala 2023; Mou et al. 2024; Ju et al. 2023b; Rombach et al. 2022) across several samples, as shown in Figure 7. We found that SD1.5 lacked pose alignment capability due to rely on solely text prompt. While T2I-Adapter and ControlNet exhibited good text semantic ability, their pose accuracy was poor, and HumanSD also showed differences in pose alignment. In contrast, our GRPose demonstrated superior pose alignment, significantly improving the details in human image generation. We also verified the model’s ability to handle multiple poses in a single condition, as shown in Figure 4. Compared to ControlNet, our model maintained accurate alignment with multiple poses, proving the feasibility of the graph learning approach. Our model addressed issues such as unnatural body positions and misaligned pose, significantly enhancing the accuracy of pose estimation. We performed transfer verification of the base model across different SD 1.5 variants from the open-source community Civitai, as shown in Figure 5. Our GRPose, when applied to different base models, enhanced their pose-guided alignment capabilities, demonstrating that GRPose is a plug-and-play solution for various diffusion models.
Figure 6: Visual comparison of PGI. Our model with PGI demonstrates enhanced pose guidance due to graph associations.
Table 2: Results of each component. $\mathcal{L}_{p}$ represents the Pose Perception Loss. These results verify the effectiveness of each component in our framework.
[…]
ControlNet 39.52 69.19 1.54
+PGI 47.80 70.63 1.51
+$\mathcal{L}_{p}$ 43.81 70.37 1.48
+PGI+$\mathcal{L}_{p}$ 49.49 70.83 1.43
Table 3: Results of pose guidance in upper body and full body on the cartoon subset from Human-Art dataset.
[…]
Upper-Body ControlNet 15.51 66.45 0.63
Ours 31.87 70.05 0.57
Full-Body ControlNet 30.78 70.12 1.29
Ours 39.19 71.03 1.17
Figure 7: Visual comparison of ours and other methods. The samples are from the Human-Art dataset, with each row representing a sample along with its corresponding pose and prompt. Our GRPose, HumanSD, ControlNet, and T2I-Adapter were provided with both a prompt and a pose, while SD1.5 was provided with only a prompt. Our model achieved outstanding results in visual quality and pose alignment.
Table 4: Results of different graph number. The use of graph convolution layers at different stages significantly improves pose alignment. This experiment represents our empirical exploration.
[…]
0 41.85 69.96 1.57
1 42.75 70.56 1.52
2 46.59 70.59 1.58
3 47.80 70.63 1.51
Table 5: Results of different loss weights $\alpha$. When $\alpha$ was set to 0.01, our model yielded the best performance.
[…]
0.1 45.58 70.36 1.52
0.01 49.49 70.83 1.43
0.05 48.44 70.68 1.45
Ablation Studies
All Components.
We conducted experiments on the Human-Art dataset to validate the effectiveness of each component, as shown in Table 2. Compared to ControlNet, the addition of the Pose Perception Loss $\mathcal{L}_{p}$ led to an increase of 4.29% in AP and 1.18% in SAP, demonstrating that the pose perception loss effectively constrained the pose differences in the feature space from an alternative perspective. Incorporating the PGI further improved performance of our model, with AP and SAP increasing by 8.28% and 1.44% respectively. This improvement indicated that PGI, through its integration of complex graph topological relationships, accurately captured high-order relationships between pose parts. It was evident that after PGI associating the information of different pose parts, our model produced significant improvements in both AP and SAP, addressing previous issues such as blurred limbs and unrealistic poses, as shown in Figure 4.
To verify the ability of generating portraits, we conducted additional experiments on upper body and full body samples from the cartoon subset on the Human-Art dataset respectively, as shown in Table 3. In the upper body scenario, our model outperformed ControlNet by 16.36% in the AP, while in the full body scenario it was 8.35% higher in the AP. The results showed that compared to ControlNet, our model had more outstanding performance in upper body generation than in full body generation, further verifying the effectiveness of our PGI associations. The performance in the upper body scenario indicated that our model achieved good results in portrait generation, which demonstrates the high potential in many downstream applications.
Graph Stages.
We conducted empirical exploration to assess the performance of our model with varying numbers of graph structures applied to the latent, pose, and fused features, as shown in Table 4. The results indicated a significant performance increase as the number of graphs increased from 1 to 4. We did not report more results due to computational resource limitations.
In addition, we visualized the change of pose alignment with and without PGI, as shown in Figure 6. In the first row of samples, the leg pose showed a noticeable error and was partially missing without PGI. In the second row, the arm was in an incorrect pose. With PGI, pose guidance significantly improved image quality, bringing the generated pose of the image closer to the origin. This demonstrates that capturing high-order relationships through the graph structure effectively facilitates information propagation between different pose parts.
Pose Perception Loss.
We examined the impact of different loss weights $\alpha$ for the Pose Perception Loss $\mathcal{L}_{p}$, as shown in Table 5. It can be found that the optimal loss weight is $0.01$. However, when the weight was set to 0.01, there was a significant decrease in AP, suggesting that it imposed constraints on the diffusion loss. A larger weight for the pose perception loss may cause the model to learn interference in the feature space.
Conclusion
In this paper, we proposed a framework, named Graph Relation Pose (GRPose), for pose-guided human image generation. Unlike to previous methods that inject pose priors into diffusion models based on Euclidean space, we used GCNs to capture the higher ordeer relationships between different pose parts. Following the ControlNet approach, our method utilizes both a text prompt and a pose prior to generate human images. We designed a Progressive Graph Integrator (PGI), which adopts a hierarchical structure within the Adapter to ensure that higher-order associations gradually correct the pose. Additionally, we proposed a novel pose perception loss, which uses a pre-trained pose estimation network to constrain the network, further enhancing the accuracy of pose guidance. Our approach achieves SOTA performance on the Human-Art and LAION-Human datasets with the best performance in the accuracy of pose alignment.